# Multi-Head RAG: Solving Multi-Aspect Problems with LLMs

### v1

@@ 4 Evaluation
Comparison Baselines We compare MRAG to two main baselines: Standard RAG and Split RAG. The first represents a modern RAG pipeline in which each document uses the activations of the last decoder layer as its embedding. The second is a blend between Standard RAG and MRAG. Specifically, it splits the activation of the last decoder layer in the same way as MRAG and applies a voting strategy. The purpose of Split RAG is to show that MRAG’s benefits come from using the multi-head output as embedding and not merely using multiple embedding spaces. Additionally, we consider Fusion RAG [30], an optional mechanism that we harness to further enhance the benefits of MRAG at the cost of additional tokens (detailed in Section 4.2).
@@
4.1 Analysis of Results
@@
Figure 4: Retrieval success ratio over 25 queries between MRAG and Standard RAG, where each query includes 10 different aspects. The upper part presents exact document matches while the lower part presents category only matches (we explain the metrics used in Section 3). A histogram is presented for a specific sample to showcase the detailed distribution among the 25 queries (the number of documents fetched for each query is 30).
@@
MRAG Standard RAG | MRAG Standard RAG | MRAG Standard RAG | MRAG Standard RAG
@@ 4.3 Benefits from Multi-Head Attention Solely
Figure 6: Relative retrieval improvements of MRAG over Standard RAG for the SFR embedding model compared with Split RAG (the blue plots), and the relative retrieval improvements of Fusion MRAG over both Fusion RAG and MRAG (the red plots).
@@ 4.5 Additional Analyses
Figure 8: Evaluation of different voting strategies for MRAG and Split RAG
We also analyze the impact of using embeddings from different decoder blocks for MRAG (instead of the last one) and the impact of using different voting strategies for MRAG as well as for Split RAG.
We consider taking multi-aspect embeddings from three different layers of the embedding model: after the first multi-head attention block, after multi-head attention block 16 (in the middle of the decoder architecture), and the final multi-head attention. We discover that the last multi-head attention performs the best when compared with the Standard RAG.
We also illustrate selected representative data from a long investigation two additional voting strategies for MRAG. We compare MRAG (1) where only the exponential lowering of significance of selected chunks is applied ($w_{i,p}=2^{-p}$), and MRAG (2) which assigns the weight for each text chunk based on the distance between the particular text chunk ($d_{i,p}$) and the query ($q$) ($w_{i}=\frac{1}{distance(d_{i,p},q)})$. Figure 8 shows that these voting strategies perform worse on average than our selected strategy for MRAG, justifying its design and selection (described in Section 2.3).
We also consider two voting strategies for Split RAG, to further deepen the empirical evaluation. Split (1) only uses the exponential lowering of significance ($w_{i,p}=2^{-p}$) and Split (2) which uses the same strategy as MRAG ($w_{i,p}=s_{i}\cdot 2^{-p}$). Figure 8 (on the right) shows that these voting strategies are on-par with each other while being worse than MRAG, further showcasing the advantages of MRAG.
@@ 5 Related Work
Many RAG schemes appeared recently [11], using the output of the last decoder layer for embedding generation. In contrast, MRAG leverages different embedding spaces of attention heads to focus on different aspects of documents and queries. As such, it can be combined with other schemes to further improve RAG pipelines.
Retrieval is sometimes enhanced by a cross-encoder reranking phase [31, 27, 28, 20, 10, 23]. In such solutions, typically after retrieving a set of relevant chunks, they are re-ranked using specialized models. In this work, we focus solely on the first retrieval phase, so MRAG can be seamlessly used in conjunction with such cross-encoders.
Structure-enhanced RAG schemes employ different strategies for structuring text to improve retrieval quality. A common idea is to construct a Knowledge Graph from text, which enables retrieval amongst entities and relationships [18, 7, 17, 4, 38]. RAPTOR [32] generates multi-level summaries for clusters of related chunks, building a tree of summaries with increasing levels of abstraction to better capture the meaning of the text. Graph RAG [8] creates a Knowledge Graph, and summarizes communities in the graph, which provide data at the different levels of abstraction. All these systems try to improve RAG quality by utilizing additional structures that describe entity relationships or the inner organization of text. Usually, they need a sophisticated preprocessing phase to prepare such structures. MRAG achieves the improvement solely based on the embedding model and has no additional storage requirements, and can be combined with any of these schemes.

### v2

@@ 4 Evaluation
Comparison Baselines We consider three main baselines: Standard RAG, Split RAG, and Fusion RAG (Rackauckas, 2024). The first represents a modern RAG pipeline in which each document uses the activations of the last decoder layer as its embedding. The second is a blend between Standard RAG and MRAG. Specifically, it splits the activation of the last decoder layer in the same way as MRAG and applies a voting strategy. The purpose of Split RAG is to show that MRAG’s benefits come from using the multi-head output as embedding and not merely using multiple embedding spaces. Additionally, we consider Fusion RAG (Rackauckas, 2024), an optional mechanism that we harness to further enhance the benefits of MRAG at the cost of additional tokens (detailed in Section 4.3).
@@
Figure 4: Retrieval success ratio over 25 queries between MRAG and Standard RAG; each query uses 10 different aspects. The top part presents exact document matches, the bottom part presents category only matches (we explain the metrics in Section 3). A histogram is presented for a specific sample to showcase the detailed distribution among the 25 queries (the number of documents fetched for each query is 30).
4.1 Superior Performance for Multi-Aspect Queries
@@
4.2 Comparable Performance for Single-Aspect Queries
@@
MRAG Standard RAG MRAG Standard RAG MRAG Standard RAG MRAG Standard RAG
@@ 4.4 Benefits from Multi-Head Attention Solely
Figure 6: Relative retrieval improvements of MRAG over Standard RAG for the SFR embedding model compared with Split RAG (the blue plots), and the relative retrieval improvements of Fusion MRAG over Standard RAG compared with Fusion RAG (the red plots).
@@ 4.6 Additional Analyses
Figure 8: Evaluation of different voting strategies for MRAG and Split RAG.
We also analyze the impact of using embeddings from different decoder blocks for MRAG (instead of the last one). Here, we consider taking multi-aspect embeddings from three different layers of the embedding model: after the first multi-head attention block, after multi-head attention block 16 (in the middle of the decoder architecture), and the final multi-head attention. We discover that the last multi-head attention performs the best when compared with the Standard RAG.
We also illustrate selected representative data from a long investigation into two additional voting strategies for MRAG. We compare MRAG (1) where only the exponential lowering of significance of selected chunks is applied ($w_{i,p}=2^{-p}$), and MRAG (2) which assigns the weight for each text chunk based on the distance between the particular text chunk ($d_{i,p}$) and the query ($q$) ($w_{i}=\frac{1}{distance(d_{i,p},q)})$. Figure 8 shows that these voting strategies perform worse on average than our selected strategy for MRAG, justifying its design and selection (described in Section 2.3). We also consider two voting strategies for Split RAG, to further deepen the empirical evaluation. Split (1) only uses the exponential lowering of significance ($w_{i,p}=2^{-p}$) and Split (2) which uses the same strategy as MRAG ($w_{i,p}=s_{i}\cdot 2^{-p}$). Figure 8 (on the right) shows that these voting strategies are on-par with each other while being worse than MRAG, further showcasing the advantages of MRAG.
The complexity of the importance score calculation (Algorithm 1) is $\mathcal{O}(n^{2})$ where $n$ is the number of the embedded documents; it is dominated by calculating the pair-wise cosine similarity and the calculation of the norm. Please note that this step needs to be done only once for each dataset and is not a bottleneck.
Finally, in addition to the performance evaluation, we also investigated the attention heads of the SFR-Embedding-Mistral model as well as LLaMA-2 7B model (model not fine-tuned for text-embedding tasks). This analysis is presented in Appendix C.
@@ 5 Related Work
Many RAG schemes appeared recently (Gao et al., 2024), using the output of the last decoder layer for embedding generation. In contrast, MRAG leverages different embedding spaces of attention heads to focus on different aspects of documents and queries. As such, it can be combined with other schemes to further improve RAG pipelines.
Retrieval is sometimes enhanced by a cross-encoder reranking phase (Rosa et al., 2022; Nogueira & Cho, 2020; Nogueira et al., 2020; Li et al., 2021; Gao et al., 2021; MacAvaney et al., 2019). In such solutions, typically after retrieving a set of relevant chunks, they are re-ranked using specialized models. In this work, we focus solely on the first retrieval phase, so MRAG can be seamlessly used in conjunction with such cross-encoders.
Structure-enhanced RAG schemes employ different strategies for structuring text to improve retrieval quality. A common idea is to construct a Knowledge Graph from text, which enables retrieval amongst entities and relationships (Jiang et al., 2024; Delile et al., 2024; Hussien et al., 2024; Bui et al., 2024; Xu et al., 2024a). RAPTOR (Sarthi et al., 2024) generates multi-level summaries for clusters of related chunks, building a tree of summaries with increasing levels of abstraction to better capture the meaning of the text. Graph RAG (Edge et al., 2024) creates a Knowledge Graph, and summarizes communities in the graph, which provide data at the different levels of abstraction. All these systems try to improve RAG quality by utilizing additional structures that describe entity relationships or the inner organization of text. Usually, they need a sophisticated preprocessing phase to prepare such structures. MRAG achieves the improvement solely based on the embedding model and has no additional storage requirements, and can be combined with any of these schemes.
@@ B.2 Dataset Details
Table 2: Overview of the structure and the number of documents in the respective datasets.
dataset #categories #topics #documents total #documents
Wikipedia 25 50 documents per category 1250
Legal Documents 25 25 per category 10 per topic 6250
Accident Reports 25 25 per category 10 per topic 6250
B.3 Prompt Template for the Synthetic Dataset Generation
@@
Appendix C Attention Head Analysis
We investigated the attention heads of two models in detail: LLaMA-2 7B and SFR-Embedding-Mistral. We selected these two models for a detailed investigation because the former represents models that are not fine-tuned for text embeddings, while the latter is specifically the text embedding model that we used for our experiments. For each model, we looked specifically at the attention scores within each attention head, i.e., how much attention each head pays to each input token during the inference. Knowing the semantics of the input tokens enables then deriving certain conclusions about multi-aspectuality and attention heads.
We plot selected results in Figure 9. Each heatmap shows the dot-product between key- and value-projections inside a given specified attention head, where line $i$ of a heatmap for attention head h indicates the dot-products between the query-projection of token $i$ and the key-projections of all previous tokens $j<i$ (both models use causal attention).
(a) Head 4 LLaMA-2
(b) Head 8 SFR
(c) Head 22 LLaMA-2
(d) Head 21 SFR
Figure 9: Heatmap plots for selected attention heads of the LLaMA-2 7B and SFR-Embedding-Mistral models.
For both models, we found out that the attention patterns vary significantly between the different attention heads. Still, we encountered two distinct patterns. First, the diagonal lines in Figures 9(a) and 9(b) indicate that, when processing a certain input token x, elevated attention is paid to some tokens that came a constant numbers of steps before x. We postulate that this pattern is likely beneficial to understanding the overall rhythm of a natural language, allowing the model to better identify which words are semantically connected, and which parts of the input text refer to each other. Second, horizontal and vertical lines in Figures 9(c) and 9(d) show that these heads learned to pay attention to specific tokens, regardless of how far apart they are within the input sequence. An intuitive justification for such patterns is the focus on certain semantic aspects of the input sequence.
(a) Attention Head 0 LLaMA-2 7B
(b) Attention Head 8 SFR
(c) Attention Head 14 LLaMA-2 7B
(d) Attention Head 21 SFR
Figure 10: Attention scores for selected attention heads of the LLaMA-2 7B and SFR-Embedding-Mistral models.
We also detail attention scores (after applying softmax) of selected heads in Figures 10 and 11, when the model is processing the last token of its input. We see that some tokens gather a lot of attention from most heads, yet there is always a plethora of passages which are attended differently by any two attention heads. An interesting pattern we encountered was that for the SFR-Embedding-Mistral model (see Figure 11), all heads’ attention spiked significantly on the first line-break in the input sequence - either positively or negatively. We conjecture that this is a consequence of how the embedding model was fine-tuned and its intended usage pattern: embedding queries are usually prepended with a retrieval instruction, which is terminated by a line-break. The model likely learnt to summarise the necessary information about this instruction inside the terminating line-break.
Figure 11: Attention scores for all attention heads of the SFR-Embedding-Mistral model.